# Estimating the size of a set using cascading exclusion

### v2

@@ 1.1 Background
There are many examples where uniform samples from $S$ are available but the size $|S|$ is unknown, and of interest. For example, if $S$ is the set of $I\times J$ contingency tables with given row and column sums, Monte Carlo Markov Chain methods give easy access to uniform samples [19] but $|S|$ is unknown. Further recent work arise using the Burnside process [38, 21, 20] for enumerating the number of orbits of a finite group acting on a finite set. This includes contingency tables as a special case [17]. More examples where uniform samples are easily accessible but $|S|$ is unknown are provided in sections 2.2, 2.3, 2.4 below.
@@ 2.1 Unseen species
$\mathbb{E}\biggl{[}\biggl{(}\frac{T_{n}}{n}-P_{n}\biggr{)}^{2}\biggr{]}\leq\frac{4}{e(n-1)}+\frac{4(n-1)}{en(n-2)}+\frac{2}{n}\leq\frac{5}{n-2}.$
@@
When $N$ is finite, one can maximize over all possible values of $p_{1},\ldots,p_{N}$ in the above theorem to obtain the following corollary of Lemma 2.2.
Lemma 2.3.
In the setting of Lemma 2.2, suppose that $N$ is finite. Then
$\mathbb{E}\biggl{[}\biggl{(}\frac{T_{n}}{n}-P_{n}\biggr{)}^{2}\biggr{]}$ $\leq\biggl{(}\frac{8}{N}+\frac{2}{n}\biggr{)}e^{-(n-2)/N}.$
Proof.
It is an easy calculus exercise to check that the upper bound in Lemma 2.2 is maximized when $p_{1}=\cdots=p_{N}=1/N$, when it is equal to
$4\sum_{i=1}^{N}\frac{1}{N^{2}}\biggl{(}1-\frac{1}{N}\biggr{)}^{n-1}+\frac{4(n-1)}{n}\sum_{i=1}^{N}\frac{1}{N^{2}}\biggl{(}1-\frac{1}{N}\biggr{)}^{n-2}$
$\qquad+\frac{2}{n}\sum_{i=1}^{N}\frac{1}{N}\biggl{(}1-\frac{1}{N}\biggr{)}^{n-1}$
$\leq\frac{4}{N}e^{-(n-1)/N}+\frac{4(n-1)}{n}\frac{1}{N}e^{-(n-2)/N}+\frac{2}{n}e^{-(n-1)/N},$
where we used the inequality $1-x\leq e^{-x}$, which holds for all $x\geq 0$. This completes the proof. ∎
@@
Take any positive integer $m$. It is easy to show that that maximum value of $x(1-x)^{m}$ as $x$ ranges over $[0,1]$ is attained when $x=1/(m+1)$, where the value is
@@
$4\sum_{i=1}^{N}\frac{p_{i}}{e(n-1)}+\frac{4(n-1)}{n}\sum_{i=1}^{N}\frac{p_{i}}{e(n-2)}+\frac{2}{n}\sum_{i=1}^{N}p_{i}\leq\frac{4}{e(n-1)}+\frac{4(n-1)}{en(n-2)}+\frac{2}{n}.$
@@
In spite of the simplicity of the bound in Theorem 2.1, the dependence on $N$ from Lemma 2.3 can be informative if $N$ is known. For example, suppose $n=N\log N-aN$ for some $a>0$. Then Lemma 2.3 shows that
$\mathbb{E}\biggl{[}\biggl{(}\frac{T_{n}}{n}-P_{n}\biggr{)}^{2}\biggr{]}$ $\leq\frac{Ce^{a}}{N^{2}},$
where $C$ is a universal constant, which is much better than $O(1/n)$ bound obtained by maximizing over $N$.
@@ 2.2 Convex hulls
$\mathbb{E}\biggl{[}\biggl{(}\frac{V_{n}}{n}-D_{n}\biggr{)}^{2}\biggr{]}\leq\frac{8d+9}{n}.$
@@
$\mathbb{E}\biggl{[}\frac{(n-V_{n})^{2}}{(8d+9)n}\biggl{(}\frac{\widehat{\mathrm{vol}(K)}}{\mathrm{vol}(K)}-1\biggr{)}^{2}\biggr{]}\leq 1.$
@@
$I_{\alpha}:=\left[\widehat{\mathrm{vol}(K)},\,\frac{\widehat{\mathrm{vol}(K)}}{1-\frac{\sqrt{(8d+9)n}}{\sqrt{\alpha}(n-V_{n})}}\right].$
To see that this indeed has level $1-\alpha$, simply note that by Markov’s inequality and the fact that $\widehat{\mathrm{vol}(K)}\leq\mathrm{vol}(K)$,
$\mathbb{P}(\mathrm{vol}(K)\notin I_{\alpha})$ $=\mathbb{P}\biggl{(}\frac{\widehat{\mathrm{vol}(K)}}{\mathrm{vol}(K)}\notin\biggl{[}1-\frac{\sqrt{(8d+9)n}}{\sqrt{\alpha}(n-V_{n})},\,1\biggr{]}\biggr{)}$
$\leq\mathbb{P}\biggl{(}\biggl{|}\frac{\widehat{\mathrm{vol}(K)}}{\mathrm{vol}(K)}-1\biggr{|}>\frac{\sqrt{(8d+9)n}}{\sqrt{\alpha}(n-V_{n})}\biggr{)}$
$=\mathbb{P}\biggl{(}\frac{(n-V_{n})^{2}}{(8d+9)n}\biggl{(}\frac{\widehat{\mathrm{vol}(K)}}{\mathrm{vol}(K)}-1\biggr{)}^{2}>\frac{1}{\alpha}\biggr{)}$
$\leq\alpha\mathbb{E}\biggl{[}\frac{(n-V_{n})^{2}}{(8d+9)n}\biggl{(}\frac{\widehat{\mathrm{vol}(K)}}{\mathrm{vol}(K)}-1\biggr{)}^{2}\biggr{]}\leq\alpha.$
@@ 2.2.1 Small illustrations in low dimension
Figure 2: Verification of the MSE bound from Theorem 2.4: $\mathbb{E}[(V_{n}/n-D_{n-1})^{2}]\leq(8d+9)/n$. The empirical mean squared error is shown for four distributions (uniform rectangle and disk, independent Gaussian, and correlated Gaussian with $r=0.8$) in dimensions $d=2$ and $d=3$. The black dashed line represents the theoretical upper bound. All empirical values fall well below the bound across different probability measures.
@@ 2.3 Interpolating between birthdays and German tanks using posets
Up-sets are a basic construct of poset theory where they are called ‘order ideals’ (see Stanley [54, Section 3.4]). A simple example of an up-set comes from using the opposite order on $S=\{1,\ldots,N\}$ and the product order on $S\times S$; the up-sets are partitions $\{(i,j),1\leq i\leq I,1\leq j\leq\lambda_{i}\mbox{ with }\lambda_{1}\geq\lambda_{2}\geq\cdots\geq\lambda_{I}\}$.
@@
$\mathbb{E}\biggl{[}\biggl{(}\frac{N_{n}}{n}-\mu(\uparrow\!\{X_{1},\ldots,X_{n}\})\biggr{)}^{2}\biggr{]}\leq\biggl{(}\frac{8}{e}+\frac{1}{2}\biggr{)}\frac{1}{n}<\frac{7}{2n}.$
@@
$\mathbb{E}\biggl{[}\frac{N_{n}^{2}}{n}\biggl{(}1-\frac{\widehat{|T|}}{|T|}\biggr{)}^{2}\biggr{]}\leq\frac{8}{e}+\frac{1}{2}.$
@@
A different motivation for this same estimor follows from the ‘capture-recapture’ work of Darroch and Ratcliff [11]. If $T$ is a finite set of size $|T|$ and $X_{1},\ldots,X_{n}$ is an i.i.d. uniform sample from $T$, let $D$ be the set of distinct values among $\{X_{1},\ldots,X_{n}\}$. The chance that the next value is not in $D$ is $\frac{|T|-|D|}{|T|}$. Estimating this using the Good–Turing estimate $\frac{n_{1}}{n}$, with $n_{1}$ the number of singletons in the sample, Darroch and Ratcliff [11] equate $\frac{|T|-|D|}{|T|}\simeq\frac{n_{1}}{n}$, which gives the estimate
@@
Note that this exactly the estimate displayed in equation (2.2). If there are many repeats, this can be improved by the Chao estimates [7, 8] that use doubletons as well as singletons.
@@ 2.4 Convex subsets of posets
$\mathbb{E}\biggl{[}\biggl{(}\frac{N_{n}}{n}-\mu(\operatorname{conv}(X_{1},\ldots,X_{n}))\biggr{)}^{2}\biggr{]}\leq\biggl{(}\frac{16}{e}+\frac{1}{2}\biggr{)}\frac{1}{n}<\frac{7}{n}.$
@@
Thus, we only have to compute upper bounds on the quantities $\theta$, $\delta^{\prime}$ and $\delta^{\prime\prime}$ from Theorem 1.1. We bound $\theta$ by $\frac{1}{4}$. Next, note that if $X_{n}\in A^{\prime}(X_{1},\ldots,X_{n-1})\Delta A^{\prime\prime}(X_{1},\ldots,X_{n-2})$, then $X_{n}\in\operatorname{conv}(X_{1},\ldots,X_{n-1})$ and $X_{n}\notin\operatorname{conv}(X_{1},\ldots,X_{n-2})$. This implies that either $X_{n}\preceq X_{n-1}$ and $X_{n}\not\preceq X_{j}$ for all $j\leq n-2$, or $X_{n-1}\preceq X_{n}$ and $X_{j}\not\preceq X_{n}$ for all $j\neq n-2$. Thus,
@@
$\mathbb{E}\biggl{[}\frac{N_{n}^{2}}{n}\biggl{(}1-\frac{\widehat{|T|}}{|T|}\biggr{)}^{2}\biggr{]}\leq\frac{16}{e}+\frac{1}{2}<7.$
@@ 2.5.2 A universal approximation theorem
$\mathbb{E}\biggl{[}\biggl{(}\frac{W_{n}(r)}{n}-\mu(U_{n}(r))\biggr{)}^{2}\biggr{]}\leq\frac{9}{n}.$
@@ 2.6 Prediction sets
$\delta_{n}^{\prime\prime}:=\mathbb{P}(Y_{n+1}\in P_{n}(X_{n+1};L_{n-1})\Delta P_{n-1}(X_{n+1};L_{n-2})).$
@@
Suppose that $(X_{1},Y_{1}),\ldots,(X_{n},Y_{n})$ that are i.i.d. from some joint distribution on $\mathbb{R}^{p}\times\mathbb{R}$. Assume that the conditional distribution of $Y_{1}$ given $X_{1}=x$ has a density with respect to Lebesgue measure that is bounded by a finite constant that does not depend on $x$. Further, assume that $Y_{1}$ and the components of $X_{1}$ have sub-Gaussian tails, and that $(X_{1},Y_{1})$ has a bounded probability density with respect to Lebesgue measure on $\mathbb{R}^{p}\times\mathbb{R}$. Let $p_{n}$ be the coverage probability of the OLS prediction interval and $\widehat{p}_{n}$ be its leave-one-out estimate. Then
@@ 2.6.1 Simulations
2.6.2 Cholostyramine data
As an application to real data, we consider an old data set from Efron and Tibshirani [27]. The response variable is an improvement score for $164$ men taking cholostyramine to reduce cholesterol. The predictor variable is their compliance. The data are available in the bootstrap package in R, and the complete R code to reproduce our analysis is available in the supplementary materials.
@@ 2.7 Connection to a theorem of Devroye and Wagner
In practice, this would be essentially be the same as the bound (2.5) from the Devroye–Wagner theorem. Still, it would be interesting to understand why there are two terms in our bound, versus only one in (2.5). In particular, would it be possible to improve Theorem 1.1 to have only one error term, corresponding to the comparison of $n$ and n1n-1? It is not immediately clear, since Theorem 1.1 has wider coverage than the Devroye–Wagner theorem. We leave this question for future investigation.
@@ Acknowledgements
We thank Louigi Addario-Berry, David Aldous, Eugenio Regazzini, Richard Stanley, and Sandy Zabell for their help. PD was supported by NSF grant DMS-1954042. SC was supported in part by NSF grants DMS-2113242 and DMS-2153654.
@@ Appendix A Appendix
$\mathbb{E}[(I_{n}-I_{n}^{\prime})^{2}]$ $=\sum_{i=1}^{n}\mathbb{E}[(K_{i}-L_{i})^{2}]+2\sum_{1\leq i<j\leq n}\mathbb{E}[(K_{i}-L_{i})(K_{j}-L_{j})].$
@@
$\mathbb{E}\biggl{|}\frac{I^{\prime}}{n}-\mu(A(X_{1},\ldots,X_{n}))\biggr{|}\leq 2\delta^{\prime}.$
@@
Let $X_{n+1}$ be a new sample drawn from $\mu$, independent of $X_{1},\ldots,X_{n}$. Let
$K_{n+1}:=1_{\{X_{n+1}\in A(X_{1},\ldots,X_{n})\}},$
so that
$\mathbb{E}(K_{n+1}|X_{1},\ldots,X_{n})=\mu(A(X_{1},\ldots,X_{n})).$ (A.1)
Next, for each $1\leq i\leq n$, define
$K_{n+1}^{i}:=1_{\{X_{n+1}\in A^{\prime}(X_{1},\ldots,X_{i-1},X_{i+1},\ldots,X_{n})\}}.$
Now, note that
$\mathbb{E}(K_{n+1}|X_{1},\ldots,X_{i-1},X_{i+1},\ldots,X_{n})$ $=\mathbb{E}(K_{i}|X_{1},\ldots,X_{i-1},X_{i+1},\ldots,X_{n})=L_{i},$
and since $K_{n+1}^{i}$ has no dependence on $X_{i}$,
$\mathbb{E}(K_{n+1}^{i}|X_{1},\ldots,X_{i-1},X_{i+1},\ldots,X_{n})=K_{n+1}^{i}.$
Thus, by the definition of $\delta^{\prime}$, we get
$\mathbb{E}|L_{i}-K_{n+1}^{i}|$ $\leq\mathbb{E}|K_{n+1}-K_{n+1}^{i}|\leq\delta^{\prime}.$
Consequently,
$\mathbb{E}|L_{i}-K_{n+1}|\leq\mathbb{E}|L_{i}-K_{n+1}^{i}|+\mathbb{E}|K_{n+1}-K_{n+1}^{i}|\leq 2\delta^{\prime}.$
Since $L_{1},\ldots,L_{n}$ have no dependence on $X_{n+1}$, this inequality, together with equation (A.1), show that
$\mathbb{E}\biggl{|}\frac{I^{\prime}}{n}-\mu(A(X_{1},\ldots,X_{n}))\biggr{|}$ $=\mathbb{E}\biggl{|}\frac{1}{n}\sum_{i=1}^{n}L_{i}-\mathbb{E}(K_{n+1}|X_{1},\ldots,X_{n})\biggr{|}$
$\leq\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}|L_{i}-K_{n+1}|\leq 2\delta^{\prime}.$
This completes the proof. ∎
@@
$\leq\frac{2\theta}{n}+\frac{4(n-1)\delta^{\prime\prime}}{n}+4\delta^{\prime}.$

### v3

# arXiv:2508.05901v3[math.ST] 27 Feb 2026

@@ 1.1 Background
There are many examples where uniform samples from $S$ are available but the size $|S|$ is unknown, and of interest. For example, if $S$ is the set of $I\times J$ contingency tables with given row and column sums, Monte Carlo Markov Chain methods give easy access to uniform samples [19] but $|S|$ is unknown. Further recent works arise using the Burnside process [38, 21, 20] for enumerating the number of orbits of a finite group acting on a finite set. This includes contingency tables as a special case [17]. More examples where uniform samples are easily accessible but $|S|$ is unknown are provided in sections 2.2, 2.3, 2.4 below.
@@ 2.1 Unseen species
$\mathbb{E}\biggl[\biggl(\frac{T_{n}}{n}-P_{n}\biggr)^{2}\biggr]\leq\frac{2}{e(n-1)}+\frac{4(n-1)}{en(n-2)}+\frac{2}{n}\leq\frac{5}{n-2}.$
@@
Take any positive integer $m$. It is easy to show that the maximum value of $x(1-x)^{m}$ as $x$ ranges over $[0,1]$ is attained when $x=1/(m+1)$, where the value is
@@
$2\sum_{i=1}^{N}\frac{p_{i}}{e(n-1)}+\frac{4(n-1)}{n}\sum_{i=1}^{N}\frac{p_{i}}{e(n-2)}+\frac{2}{n}\sum_{i=1}^{N}p_{i}\leq\frac{2}{e(n-1)}+\frac{4(n-1)}{en(n-2)}+\frac{2}{n}.$
@@
In spite of the simplicity of the bound in Theorem 2.1, the bound can be improved if $N\ll n$, as shown by the following lemma (which is a corollary of Lemma 2.2).
Lemma 2.3.
In the setting of Lemma 2.2, we have
$\mathbb{E}\biggl[\biggl(\frac{T_{n}}{n}-P_{n}\biggr)^{2}\biggr]$ $\leq\frac{4N}{(n-2)^{2}}$
Proof.
Using the bound $1-x\leq e^{-x}$, we get that the bound in Lemma 2.2 is bounded above by $\sum_{i=1}^{N}g(p_{i})$, where
$g(p)$ $:=2p^{2}(1-p)e^{-(n-2)p}+\frac{4(n-1)}{n}p^{2}e^{-(n-2)p}+\frac{2}{n}p(1-p)e^{-(n-2)p}$
$\leq\biggl(6p^{2}+\frac{2p}{n}\biggr)e^{-(n-2)p}.$
Writing t:=(n2)pt:=(n-2)p, the last expression can be expressed as
$\biggl(\frac{6t^{2}}{(n-2)^{2}}+\frac{2t}{n(n-2)}\biggr)e^{-t}\leq\frac{(6t^{2}+2t)e^{-t}}{(n-2)^{2}}.$
Now, $t^{2}e^{-t}$ is maximized at $t=2$ and $te^{-t}$ is maximized at $t=1$. Thus, for any $t>0$,
$(6t^{2}+2t)e^{-t}\leq 24e^{-2}+2e^{-1}\leq 4.$
This gives the required bound. ∎
@@ 2.2 Convex hulls
$\mathbb{E}\biggl[\biggl(\frac{V_{n}}{n}-D_{n}\biggr)^{2}\biggr]\leq\frac{6d+7}{n}.$
@@
$\mathbb{E}\biggl[\frac{(n-V_{n})^{2}}{(6d+7)n}\biggl(\frac{\widehat{\mathrm{vol}(K)}}{\mathrm{vol}(K)}-1\biggr)^{2}\biggr]\leq 1.$
@@
$I_{\alpha}:=\left[\frac{\widehat{\mathrm{vol}(K)}}{1+\frac{\sqrt{(6d+7)n}}{\sqrt{\alpha}(n-V_{n})}},\,\frac{\widehat{\mathrm{vol}(K)}}{\max\biggl\{1-\frac{\sqrt{(6d+7)n}}{\sqrt{\alpha}(n-V_{n})},0\biggr\}}\right].$
To see that this indeed has level $1-\alpha$, simply note that by Markov’s inequality,
$\mathbb{P}(\mathrm{vol}(K)\notin I_{\alpha})$ $=\mathbb{P}\biggl(\frac{\widehat{\mathrm{vol}(K)}}{\mathrm{vol}(K)}\notin\biggl[1-\frac{\sqrt{(6d+7)n}}{\sqrt{\alpha}(n-V_{n})},\,1+\frac{\sqrt{(6d+7)n}}{\sqrt{\alpha}(n-V_{n})}\biggr]\biggr)$
$=\mathbb{P}\biggl(\biggl|\frac{\widehat{\mathrm{vol}(K)}}{\mathrm{vol}(K)}-1\biggr|>\frac{\sqrt{(6d+7)n}}{\sqrt{\alpha}(n-V_{n})}\biggr)$
$=\mathbb{P}\biggl(\frac{(n-V_{n})^{2}}{(6d+7)n}\biggl(\frac{\widehat{\mathrm{vol}(K)}}{\mathrm{vol}(K)}-1\biggr)^{2}>\frac{1}{\alpha}\biggr)$
$\leq\alpha\mathbb{E}\biggl[\frac{(n-V_{n})^{2}}{(6d+7)n}\biggl(\frac{\widehat{\mathrm{vol}(K)}}{\mathrm{vol}(K)}-1\biggr)^{2}\biggr]\leq\alpha.$
@@ 2.2.1 Small illustrations in low dimension
Figure 2: Verification of the MSE bound from Theorem 2.4: $\mathbb{E}[(V_{n}/n-D_{n})^{2}]\leq(6d+7)/n$. The empirical mean squared error is shown for four distributions (uniform rectangle and disk, independent Gaussian, and correlated Gaussian with $r=0.8$) in dimensions $d=2$ and $d=3$. The black dashed line represents the theoretical upper bound. All empirical values fall well below the bound across different probability measures.
@@ 2.3 Interpolating between birthdays and German tanks using posets
Up-sets are a basic construct of poset theory where they are called ‘order filters’ (see Stanley [54, Section 3.4]). A simple example of an up-set comes from using the opposite order on $S=\{1,\ldots,N\}$ and the product order on $S\times S$; the up-sets are partitions $\{(i,j),1\leq i\leq I,1\leq j\leq\lambda_{i}\mbox{ with }\lambda_{1}\geq\lambda_{2}\geq\cdots\geq\lambda_{I}\}$.
@@
$\mathbb{E}\biggl[\biggl(\frac{N_{n}}{n}-\mu(\uparrow\!\{X_{1},\ldots,X_{n}\})\biggr)^{2}\biggr]\leq\biggl(\frac{6}{e}+\frac{1}{2}\biggr)\frac{1}{n}<\frac{3}{n}.$
@@
$\mathbb{E}\biggl[\frac{N_{n}^{2}}{n}\biggl(1-\frac{\widehat{|T|}}{|T|}\biggr)^{2}\biggr]\leq\frac{6}{e}+\frac{1}{2}.$
@@
A different motivation for this same estimator follows from the ‘capture-recapture’ work of Darroch and Ratcliff [11]. If $T$ is a finite set of size $|T|$ and $X_{1},\ldots,X_{n}$ is an i.i.d. uniform sample from $T$, let $D$ be the set of distinct values among $\{X_{1},\ldots,X_{n}\}$. The chance that the next value is not in $D$ is $\frac{|T|-|D|}{|T|}$. Estimating this using the Good–Turing estimate $\frac{n_{1}}{n}$, with $n_{1}$ the number of singletons in the sample, Darroch and Ratcliff [11] equate $\frac{|T|-|D|}{|T|}\simeq\frac{n_{1}}{n}$, which gives the estimate
@@
Note that this is exactly the estimate displayed in equation (2.2). If there are many repeats, this can be improved by the Chao estimates [7, 8] that use doubletons as well as singletons.
@@ 2.4 Convex subsets of posets
$\mathbb{E}\biggl[\biggl(\frac{N_{n}}{n}-\mu(\operatorname{conv}(X_{1},\ldots,X_{n}))\biggr)^{2}\biggr]\leq\biggl(\frac{12}{e}+\frac{1}{2}\biggr)\frac{1}{n}<\frac{5}{n}.$
@@
Thus, we only have to compute upper bounds on the quantities $\theta$, $\delta^{\prime}$ and $\delta^{\prime\prime}$ from Theorem 1.1. We bound $\theta$ by $\frac{1}{4}$. Next, note that if $X_{n}\in A^{\prime}(X_{1},\ldots,X_{n-1})\Delta A^{\prime\prime}(X_{1},\ldots,X_{n-2})$, then $X_{n}\in\operatorname{conv}(X_{1},\ldots,X_{n-1})$ and $X_{n}\notin\operatorname{conv}(X_{1},\ldots,X_{n-2})$. This implies that either $X_{n}\preceq X_{n-1}$ and $X_{n}\not\preceq X_{j}$ for all $j\leq n-2$, or $X_{n-1}\preceq X_{n}$ and $X_{j}\not\preceq X_{n}$ for all $j\leq n-2$. Thus,
@@
$\mathbb{E}\biggl[\frac{N_{n}^{2}}{n}\biggl(1-\frac{\widehat{|T|}}{|T|}\biggr)^{2}\biggr]\leq\frac{12}{e}+\frac{1}{2}<5.$
@@ 2.5.2 A universal approximation theorem
$\mathbb{E}\biggl[\biggl(\frac{W_{n}(r)}{n}-\mu(U_{n}(r))\biggr)^{2}\biggr]\leq\frac{7}{n}.$
@@ 2.6 Prediction sets
$\delta_{n}^{\prime\prime}:=\mathbb{P}(Y_{n+1}\in P_{n-1}(X_{n+1};L_{n-1})\Delta P_{n-2}(X_{n+1};L_{n-2})).$
@@
Suppose that $(X_{1},Y_{1}),\ldots,(X_{n},Y_{n})$ are i.i.d. from some joint distribution on $\mathbb{R}^{p}\times\mathbb{R}$. Assume that the conditional distribution of $Y_{1}$ given $X_{1}=x$ has a density with respect to Lebesgue measure that is bounded by a finite constant that does not depend on $x$. Further, assume that $Y_{1}$ and the components of $X_{1}$ have sub-Gaussian tails, and that $(X_{1},Y_{1})$ has a bounded probability density with respect to Lebesgue measure on $\mathbb{R}^{p}\times\mathbb{R}$. Let $p_{n}$ be the coverage probability of the OLS prediction interval and $\widehat{p}_{n}$ be its leave-one-out estimate. Then
@@ 2.6.1 Simulations
2.6.2 Cholestyramine data
As an application to real data, we consider an old data set from Efron and Tibshirani [27]. The response variable is an improvement score for $164$ men taking cholestyramine to reduce cholesterol. The predictor variable is their compliance. The data are available in the bootstrap package in R, and the complete R code to reproduce our analysis is available in the supplementary materials.
@@ 2.7 Connection to a theorem of Devroye and Wagner
In practice, this would be essentially the same as the bound (2.5) from the Devroye–Wagner theorem. Still, it would be interesting to understand why there are two terms in our bound, versus only one in (2.5). In particular, would it be possible to improve Theorem 1.1 to have only one error term, corresponding to the comparison of $n$ and n1n-1? It is not immediately clear, since Theorem 1.1 has wider coverage than the Devroye–Wagner theorem. We leave this question for future investigation.
@@ Acknowledgements
We thank Louigi Addario-Berry, David Aldous, Eugenio Regazzini, Richard Stanley, Tim Sudijono, and Sandy Zabell for their help. PD was supported by NSF grant DMS-1954042. SC was supported in part by NSF grants DMS-2113242 and DMS-2153654.
@@ Appendix A Appendix
$\mathbb{E}[(I-I^{\prime})^{2}]$ $=\sum_{i=1}^{n}\mathbb{E}[(K_{i}-L_{i})^{2}]+2\sum_{1\leq i<j\leq n}\mathbb{E}[(K_{i}-L_{i})(K_{j}-L_{j})].$
@@
$\mathbb{E}\biggl|\frac{I^{\prime}}{n}-\mu(A(X_{1},\ldots,X_{n}))\biggr|\leq\delta^{\prime}.$
@@
Note that by definition of $L_{i}$ and independence of the $X_{j}$’s,
$L_{i}$ $=\mu(A^{\prime}(X_{1},\ldots,X_{i-1},X_{i+1},\ldots,X_{n})).$
Thus,
$\frac{I^{\prime}}{n}$ $=\frac{1}{n}\sum_{i=1}^{n}\mu(A^{\prime}(X_{1},\ldots,X_{i-1},X_{i+1},\ldots,X_{n})).$
By the inequality $|\mu(B)-\mu(C)|\leq\mu(B\Delta C)$, this gives
$\biggl|\frac{I^{\prime}}{n}-\mu(A(X_{1},\ldots,X_{n}))\biggr|$ $\leq\frac{1}{n}\sum_{i=1}^{n}\mu(A^{\prime}(X_{1},\ldots,X_{i-1},X_{i+1},\ldots,X_{n})\Delta A(X_{1},\ldots,X_{n})).$
Taking expectation on both sides and using symmetry, we get the desired result. ∎
@@
$\leq\frac{2\theta}{n}+\frac{4(n-1)\delta^{\prime\prime}}{n}+2\delta^{\prime}.$